OpenStudy (kanwal32):

tan^3 theta + cot^3 theta=52, then the value of tan^2theta is equal to A)14 B)15 C)16 D)17

OpenStudy (kanwal32):

Answer is ABD but how \[\sin ^{6} + \cos ^{6}/\sin ^{3}\cos ^{3}\]

OpenStudy (kanwal32):

I have simplified in this form but what to do next should I put values and check

OpenStudy (kanwal32):

pls can anyone answer day after tomorrow is my exam